Calculate the number of grams of water that must be added to 10.0 g of phenol to prepare a 2.00 percent aqueous phenol solution.


1
Expert's answer
2021-04-06T06:07:48-0400

"\frac{10×100}{x}" = 2


"x = 500g"

mass of solution is 500g

mass of water = 500-10 = 490g

490 g of water must be added